OpenStudy (anonymous):

how do you find the period of the graph of an equation, general answer, if needed, here is an equation y=2sin1/3x

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

just clarifying... is it \(\large y=2sin(\frac{1}{3}x) \) ????

OpenStudy (anonymous):

just like that, though without the parenthesis

OpenStudy (anonymous):

i've heard there is a formula used, not too sure

OpenStudy (anonymous):

the function is written in the form y = A*sin(Bx + C) + D, so the period is \(\large \frac{2\pi}{B} \) where B =\(\large \frac{1}{3} \) in your equation.

OpenStudy (anonymous):

so the answer is 6(pi) am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)